Octavia Bakery is planning to purchase one of two ovens. The expected cash flows for each oven are shown below. MARR is 10%/What is the discounted payback period for Model 334A?

Entry field with incorrect answer
Answer #1

For finding a discounted payback period we need to find the present value of future revenue and then find Cumulative cash flow (CCF), we need to find the year in which CCF turns positive

We will use formula P = F/(1+i)^t to find the present value of future cash flows

Discounted payback period = Year bef CCF turns positive + (Absolute value of CCF bef it turns positive/Present value of cash flow in the year in which CCF turns positive)

using excel

Year Investment Annual revenue Annual Cost Salvage value Net cash flow Present value of Net Cash Flow CCF
0 -74000 -74000 -74000.00 -74000
1 28500 -6200 22300 20272.73 -53727.27
2 28500 -6200 22300 18429.75 -35297.52
3 28500 -6200 22300 16754.32 -18543.20
4 28500 -6200 22300 15231.20 -3312.00
5 28500 -6200 0 22300 13846.55 10534.54

Discounted payback period = 4 + 3312 / 13846.55

= 4 + 0.24

= 4.24 yrs

Showing formula in excel

Year Investment Annual revenue Annual Cost Salvage value Net cash flow Present value of Net Cash Flow CCF
0 -74000 =B153+C153+D153+E153 =F153/(1+0.1)^A153 =G153
1 28500 -6200 =B154+C154+D154+E154 =F154/(1+0.1)^A154 =H153+G154
2 28500 -6200 =B155+C155+D155+E155 =F155/(1+0.1)^A155 =H154+G155
3 28500 -6200 =B156+C156+D156+E156 =F156/(1+0.1)^A156 =H155+G156
4 28500 -6200 =B157+C157+D157+E157 =F157/(1+0.1)^A157 =H156+G157
5 28500 -6200 0 =B158+C158+D158+E158 =F158/(1+0.1)^A158 =H157+G158